OpenStudy (anonymous):

Use the given domain to find the range y=-2x+3 D: x is less then or equal to 0

OpenStudy (anonymous):

Input interpretation: f(x)=-2x+3 over the interval [0, infinity) Lets say x=0, (and this is the largest value for -2x that we can get, since the rest of the x-values make it negative) f(0)=-2(0)+3 f(0)= 3 So the greatest range is 3. And the range approaches infinity closer and closer, the larger the x (the domain) gets.

OpenStudy (anonymous):

Hope I helped.